An AM radio station 20km away from you is broadcasting at a frequency of 1000 kHz. Approximately how many waves are there between you and the station?
Answer #1

Start with c = f(wavelength)

3 X 108 = 1000000(wavelength)

wavelength = 300 m

Then total distance = 20000 m

20000/300 = 66.7 waves

So round up to 67 waves for the approximate
Answer #2

Wavelength = Speed of wave/Frequency

= 3*10^8/1000*10^3

= 300 Meter

Number of waves in between 20 Km way = 20,000/300 = 66.67 ~67 Waves